2. Find the derivative and simplify. y = (3x-4)2 (1-x²)6
Answer #1

Given that Y= (37-48x (1-22 -6 Differentiating dy = 2 (34-u) x(3-0)*(1-2936+(-6)*(1-)} (0-22)X Wich. to x dai (31-422 (h-ko= 6(20-W xh 1-9246a²-8a 1-H2 = 6(3X-4) x *{!? 1+572-842 (1-x296 1-42 = 6 (32-WX 4542-84+1} (1-927 ..dy da = 6039-4){582-89 +1